# Relative-error monotonicity testing

### v1

@@ 7.2 Proof of Lemma 33
Now we start the proof of Lemma 33. Formally, a deterministic, adaptive, $q$-query algorithm ALG is a depth-$q$ binary decision tree, in which each internal node $u$ is labelled with a query point $x_{u}$ and has two children with edges to them labelled with 00 and $1$, respectively. Each leaf of the tree is labelled either “accept” or “reject.” Given any function $f$, the execution of ALG on $f$ induces a path in the tree and ALG returns the label of the leaf at the end. Since the goal of ALG is to distinguish $\mathcal{D}_{\mathrm{yes}}$ from $\mathcal{D}_{\mathrm{no}}$ (as described in the statement of Lemma 33), we may assume without loss of generality that every point queried in ALG lies in the two layers of $43n/$ 4 and $13n/4+$ 1.
@@
$\sum_{i\in[L]}|P_{i}|\leq 2|Q|\quad\text{and}\quad\sum_{j\in[M]}|P_{i,j}|\leq 2|P_{i}|,\quad\text{for every $i\in[L]$.}$
@@
$\mathop{{\bf Pr}\/}_{({\bm{T}},\mathbf{C})\sim\mathcal{E},{\bm{H}}\sim\mathcal{E}_{\text{yes}}}\big{[}\text{outcome of $\text{ALG}$ on $f_{{\bm{T}},\mathbf{C},{\bm{H}}}$ is $\ell$}\big{]}$
$\hskip 113.81102pt\leq(1+\alpha)\cdot\mathop{{\bf Pr}\/}_{({\bm{T}},\mathbf{C})\sim\mathcal{E},{\bm{H}}\sim\mathcal{E}_{\text{no}}}\big{[}\text{outcome of $\text{ALG}$ on $f_{{\bm{T}},\mathbf{C},{\bm{H}}}$ is $\ell$}\big{]}.$
@@
The proof is similar to that of 28.
@@
$\dfrac{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{no}}}[(T,C,{\bm{H}})\text{ results in outcome $\ell$}]}{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{yes}}}[(T,C,{\bm{H}})\text{ results in outcome $\ell$}]}\geq\frac{1}{1+\alpha}.$ (5)
Let $I$ be the set of $i\in[L]$ such that $P_{i}\neq\emptyset$; for each $i\in[L]$, let $J_{i}$ be the set of $j\in[M]$ such that $P_{i,j}\neq\emptyset$. Using (4) and an argument similar to that used in 28, this ratio is at least
@@
We prove 37 in Section 7.3. Lemma 33 follows from 36 and 37, with a proof similar to that of Lemma 23.
7.3 Proof of 37
@@
We prove the following two claims, from which 37 follows. We prove 39 first, and then prove 38 in the rest of the section.
@@
The outcome of $\text{ALG}^{\prime}$ on ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}$ satisfies Condition $C_{1}$ with probability $1-o(1)$.
@@
$\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}}\big{[}\text{outcome of $\text{ALG}$ on ${\bm{f}}$ is $\ell$}\big{]}\geq(1-\alpha/2)\cdot\mathop{{\bf Pr}\/}_{{\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}}\big{[}\text{outcome of $\text{ALG}^{\prime}$ on ${\bm{f}}$ is $\ell$}\big{]}.$
@@
$\frac{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{D}_{\mathrm{yes}}}[\text{outcome of $\text{ALG}$ on $f_{T,C,{\bm{H}}}$ is $\ell$}]}{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{D}_{\mathrm{yes}}}[\text{outcome of $\text{ALG}^{\prime}$ on $f_{T,C,{\bm{H}}}$ is $\ell$}]}\geq 1-\frac{\alpha}{2}.$
@@
again using Condition $C_{1}$. The rest of the proof follows from calculations similar to those at the end of the proof of 36. ∎
To prove 38, we start with a sufficient condition for the outcome of $\text{ALG}^{\prime}$ to satisfy Condition $C_{1}$. We say the execution of $\text{ALG}^{\prime}$ on $f$ from the support of $\mathcal{D}_{\mathrm{yes}}$ is regular if the following condition holds in every round: Let $(Q;P;R;\rho)$ be the outcome of queries revealed to $\text{ALG}^{\prime}$ so far (recall the modification of $\rho$ since we are interested in $\text{ALG}^{\prime}$ instead of ALG), let $x$ be the next query, and let $(Q^{\prime};P^{\prime};R^{\prime};\rho^{\prime})$ be the updated outcome after $x$ is queried. Then they satisfy
@@
with $A_{i,j,1}$ being defined similarly (using $y_{k}=1$ instead of 00).
@@
The execution of $\text{ALG}^{\prime}$ on ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}$ is regular with probability at least $1-o(1)$.
@@ Appendix A Separating standard and relative-error testing
It is clear that every function in the support of ${\cal D}_{\mathrm{yes}}$ belongs to ${\cal C}$, whereas every function in the support of ${\cal D}_{\mathrm{no}}$ has $\mathsf{rel}\text{-}\mathsf{dist}(f,{\cal C})=1>\varepsilon$. We claim that any relative-error $\varepsilon$-testing algorithm for ${\cal C}$ must make $\Omega(2^{n/3})$ black-box queries or samples. To see this, first observe that any algorithm making $o(2^{n/3})$ black-box queries has only an $o(1)$ chance of every querying an input $x$ such that $\bm{f}(x)=1$, whether $\bm{f}$ is drawn from ${\cal D}_{\mathrm{yes}}$ or from ${\cal D}_{\mathrm{no}}$, and thus the two distributions are indistinguishable from $o(2^{n/3})$ queries. But it is also the case that any algorithm making $m=o(2^{n/3})$ calls to $\mathrm{Samp}(\bm{f})$ will with probability $1-o(1)$ receive an identically distributed sample in both cases ($\bm{f}\sim{\cal D}_{\mathrm{yes}}$ versus $\bm{f}\sim{\cal D}_{\mathrm{no}}$), since in both cases the sample will consist of $m$ independent uniform distinct elements of $\{0,1\}^{n}$.
@@ Appendix B Both oracles are needed for relative-error monotonicity testing
    Let $\bm{S}$ be a uniform random subset of $2^{0.47n}$ of the (${n\choose n/2}=\Theta(2^{n}/\sqrt{n})$ many strings in $\{0,1\}^{n}$ that have exactly $n/2$ ones. The function $f$ outputs 1 on input $x$ if either $x_{1}+\cdots+x_{n}\geq 9n/10$ or if $x\in\bm{S}$.
In case (2), a simple argument shows that any $2^{n/2}$-query algorithm such as $A$ will only have an $o(1)$ probability of querying an input (an element of $\bm{S}$) on which case (1) and case (2) disagree, so with $1-o(1)$ probability the sequence of responses to queries that $A$ receives will be identical in these two cases. Since in case (1) the function $f$ is monotone and in case (2) $f$ has relative-distance at least 0.49 from every monotone function, such an algorithm $A$ cannot be a successful relative-error monotonicity tester.

### v2

@@ 6.2 Proof of Lemma 35
Now we start the proof of Lemma 35. Formally, a deterministic, adaptive, $q$-query algorithm ALG is a depth-$q$ binary decision tree, in which each internal node $u$ is labelled with a query point $x_{u}$ and has two children with edges to them labelled with 0 and $1$, respectively. Each leaf of the tree is labelled either “accept” or “reject.” Given any function $f$, the execution of ALG on $f$ induces a path in the tree and ALG returns the label of the leaf at the end. Since the goal of ALG is to distinguish $\mathcal{D}_{\mathrm{yes}}$ from $\mathcal{D}_{\mathrm{no}}$ (as described in the statement of Lemma 35), we may assume without loss of generality that every point queried in ALG lies in the two layers of $3n/4$ and $3n/4+1$.
@@
[…]
@@
[…]
@@
The proof is similar to that of 27.
@@
[…]
Let $I$ be the set of $i\in[L]$ such that $P_{i}\neq\emptyset$; for each $i\in[L]$, let $J_{i}$ be the set of $j\in[M]$ such that $P_{i,j}\neq\emptyset$. Using (11) and an argument similar to that used in 27, this ratio is at least
@@
We prove 39 in Section 6.3. Lemma 35 follows from 38 and 39, with a proof similar to that of Lemma 21.
6.3 Proof of 39
@@
We prove the following two claims, from which 39 follows. We prove 41 first, and then prove 40 in the rest of the section.
@@
The outcome of $\text{ALG}^{\prime}$ on ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}$ satisfies Condition $C_{1}$ with probability 1o(1)1-o(1).
@@
[…]
@@
[…]
@@
again using Condition $C_{1}$. The rest of the proof follows from calculations similar to those at the end of the proof of 38. ∎
To prove 40, we start with a sufficient condition for the outcome of $\text{ALG}^{\prime}$ to satisfy Condition $C_{1}$. We say the execution of $\text{ALG}^{\prime}$ on $f$ from the support of $\mathcal{D}_{\mathrm{yes}}$ is regular if the following condition holds in every round: Let $(Q;P;R;\rho)$ be the outcome of queries revealed to $\text{ALG}^{\prime}$ so far (recall the modification of $\rho$ since we are interested in $\text{ALG}^{\prime}$ instead of ALG), let $x$ be the next query, and let $(Q^{\prime};P^{\prime};R^{\prime};\rho^{\prime})$ be the updated outcome after $x$ is queried. Then they satisfy
@@
with $A_{i,j,1}$ being defined similarly (using $y_{k}=1$ instead of 0).
@@
The execution of $\text{ALG}^{\prime}$ on ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}$ is regular with probability at least 1o(1)1-o(1).
@@ Appendix A Separating standard and relative-error testing
It is clear that every function in the support of ${\cal D}_{\mathrm{yes}}$ belongs to $\cal C$}, whereas every function in the support of ${\cal D}_{\mathrm{no}}$ has $\mathsf{rel}\text{-}\mathsf{dist}(f,{\cal C})=1>\varepsilon$. We claim that any relative-error $\varepsilon$-testing algorithm for $\cal C$} must make $\Omega(2^{n/3})$ black-box queries or samples. To see this, first observe that any algorithm making $o(2^{n/3})$ black-box queries has only an $o(1)$ chance of every querying an input $x$ such that $\bm{f}(x)=1$, whether $\bm{f}$ is drawn from ${\cal D}_{\mathrm{yes}}$ or from ${\cal D}_{\mathrm{no}}$, and thus the two distributions are indistinguishable from $o(2^{n/3})$ queries. But it is also the case that any algorithm making $m=o(2^{n/3})$ calls to $\mathrm{Samp}(\bm{f})$ will with probability 1o(1)1-o(1) receive an identically distributed sample in both cases ($\bm{f}\sim{\cal D}_{\mathrm{yes}}$ versus $\bm{f}\sim{\cal D}_{\mathrm{no}}$), since in both cases the sample will consist of $m$ independent uniform distinct elements of $\{0,1\}^{n}$.
@@ Appendix B Both oracles are needed for relative-error monotonicity testing
    Let $\bm{S}$ be a uniform random subset of $2^{0.47n}$ of the ${n\choose n/2}=\Theta(2^{n}/\sqrt{n})$ many strings in $\{0,1\}^{n}$ that have exactly $n/2$ ones. The function $f$ outputs 1 on input $x$ if either $x_{1}+\cdots+x_{n}\geq 9n/10$ or if $x\in\bm{S}$.
In case (2), a simple argument shows that any $2^{n/2}$-query algorithm such as $A$ will only have an $o(1)$ probability of querying an input (an element of $\bm{S}$) on which case (1) and case (2) disagree, so with 1o(1)1-o(1) probability the sequence of responses to queries that $A$ receives will be identical in these two cases. Since in case (1) the function $f$ is monotone and in case (2) $f$ has relative-distance at least 0.49 from every monotone function, such an algorithm $A$ cannot be a successful relative-error monotonicity tester.